# New axisymmetric equilibria with flow from

### v1

@@ 4 Equilibrium configurations
By employing the solution to (7) derived in the previous section we have constructed several new equilibria. Particularly, we examined how the generalized Solovév equilibrium is modified by varying the additional parameters $\alpha$, $\beta$ and $\gamma$, associated with the respective linear-in-$u$ terms in (2) of pressure, poloidal-current in conjunction with the poloidal velocity, and electric field. To impose a stagnation point, $(x_{0},y_{0})$, in addition to the conditions $dU/dx|_{x_{0},y_{0}}$ and $dU/dy|_{x_{0},y_{0}}$ we considered the sign of the quantity
@@
which should be positive for a magnetic axis and negative for an X-point. For the up-down symmetric configurations considered, the magnetic axis is located at the same position, $(x=1,y=0)$ as the generalized Solovév configuration and $U=U_{a}=0$ thereon. Note that in this case the condition $dU/dy|_{(1,0)}$ is identically satisfied because of symmetry. If not otherwise specified, we assigned the following parametric values: $k=0.1$, $\delta=0.1$, $\lambda=0.1$, $D_{1}=1$ and $D_{2}=0$. It is recalled that when $\beta\neq 0$ and $\epsilon\neq 0$, $\delta$ should be calculated from (32). The parameters $a_{0}$, $b_{0}$ and $d_{0}$ were determined as solutions to the set of the algebraic equations derived by the conditions associated with the imposition of a stagnation point. Since we employed dimensionless quantities, the reference dimensional parameters $R_{0}$, $P_{s0}$ (in units of $4\pi\times 10^{-7}$ $H/m\,Pa$ ) and $U_{0}$ remained free. Depending on the value of $\epsilon$ we constructed tokamak, spherical tokamak and spheromak equilibria, the latter corresponding to $\epsilon=0$.
@@
First, we examined the impact of the $\alpha$- $\beta$- and $\gamma$-term individually on up-down symmetric equilibria equilibria by imposing the same position of the magnetic axis, $(x=1,y=0)$, as that of the respective generalized Solovév equilibrium and the same value $U=0$ thereon. In the presence of $\alpha$-term, a D-shaped configuration is formed with negative triangularity, the boundary of which is shown in Fig. 2-left by the red-dashed curve. As $\alpha$ gradually takes larger values the configuration increasing triangularity extends towards the axis of symmetry, touches the axis of symmetry forming a separatrix with a single X-point located on the origin of the coordinate system (blue, dot-dashed line) and then becomes a peculiar configuration having a separatrix with a couple of X-points (green-dotted line).
@@
The $\beta$-term creates configurations with small triangularity elongated along the $x$-axis as shown in Fig. 3. The red-dashed curves represent the magnetic surfaces of a configuration with the same poloidal magnetic flux as the respective generalized Solovév tokamak configuration represented by the black-continuous curve. As $\beta$ takes larger values the configuration gets more extended and eventually reaches the axis of symmetry becoming a spheromak one. It is noted that for $\epsilon=\beta=\gamma=0$ the equilibrium becomes independent of $\beta$. The impact of the $\gamma$-term is similar to that of the $\alpha$-term as shown in Fig. 4.
@@
Subsequently, the impact of combinations of the $\alpha$-, $\beta$- and $\gamma$-terms was examined by imposing an X-point at several positions. Thus, a variety of configurations were created. Three examples are given in Figs. 5-7. In Fig. 5 the blue, dot-dashed curves represent an up-down symmetric spherical tokamak configuration ($\epsilon=0.1$) the X-points of which are common with those of the respective generalized Solovév equilibrium, but the poloidal magnetic flux of this former equilibrium is larger than that of the latter. The red-dashed curve represents the boundary of a D-shaped configuration with poloidal magnetic flux equal to that of the generalized Solovév solution. In Fig. 6 is shown an up-down symmetric diverted configuration with the imposed X-points located at ($x=0.84,y=\pm 14.5$), while Fig. 7 shows an up-down asymmetric configuration with the magnetic axis located at ($x=1.02,y=0.83$) and the single lower X-point at $(x=0.94,y=-22.6)$.
@@
Figure 3: The impact of the additional linear-in-$U$, poloidal-current term in the second of (2) in connection with different values of $\beta$. The black continuous curves represent the magnetic surfaces of the respective generalized Solovév tokamak configuration.
@@
Figure 4: Left: different equilibrium boundaries are given as the additional linear-in-$U$, electric-field term in the third of (2) varies in connection with different values of $\gamma$. The black continuous curve represents the separatrix of the respective generalized Solovév spheromak configuration. Right: A set of magnetic surfaces for one of the configuration represented by green-dotted U-curves are shown together with the respective curves of the Solovév equilibrium.
@@
Figure 5: A spherical tokamak diverted configuration ($\epsilon=0.1$, $\delta=20$) with a couple of X-points common with the respective Solovév equilibrium. The red-dashed curve represents the boundary of a D-shaped tokamak equilibrium with poloidal magnetic flux equal to that of the Solovév one.
@@
Figure 6: An up-down symmetric, diverted tokamak configuration ($\epsilon=0.1,\ \delta=20$) with a couple of X-points located at ($x=0.84,y=\pm 14.5$). The black continuous curves represent the magnetic surfaces of the respective Solovév equilibrium.
@@
Figure 7: An up-down asymmetric, diverted tokamak configuration for $k=0.2$, $\epsilon=0.2$, $\delta=31.6$, $D_{1}=0.5$ and $D_{2}=0.01$. The magnetic axis is located at $(x=1.02,y=0.83)$ and the single lower X-point at $(x=0.94,y=-22.6)$. The black continuous curves represent the magnetic surfaces of the respective Solovév equilibrium.
@@ 5 Conclusions
Making a generic linearizing choice of the free function-terms involved in the generalized GS equation (1) we solved the resulting linearized equation (2) analytically. The requested solution was expressed as an expansion of the generalized Solovév solution (5) in the form (8) involving the determinable function $W$. Then, the solution was obtained as a superposition of the counterpart homogeneous equation and a particular solution of the complete inhomogeneous equation in the form (20). The solutions of the pertinent radial ODEs were expressed in terms of infinite converging series.
Employing these solutions, we constructed several configurations by changing the values of the free parameters, either individually or in combination, associated with the new terms of pressure, poloidal current and electric field. The equilbria constructed are pertinent to tokamaks, spherical tokamaks and spheromaks; they include D-shaped configurations with either positive or negative triangularity and diverted configurations with either a couple of X-points or a single X-point.
@@
This work was conducted in the framework of participation of the University of Ioannina in the National Programme for the Controlled Thermonuclear Fusion, Hellenic Republic.

### v2

@@ 4 Equilibrium configurations
By employing the solution to (14) derived in section 3 we have constructed several new equilibria. Particularly, we examined how the generalized Solov’ev equilibrium is modified by varying the additional parameters $\alpha$, $\beta$ and $\gamma$, associated with the respective linear-in-$u$ terms in (2) of pressure, poloidal-current in conjunction with the parallel velocity (Eqs. (5) and (7)), and electric field. To impose a stagnation point, $(x_{0},y_{0})$, in addition to the conditions $dU/dx|_{x_{0},y_{0}}=0$ and $dU/dy|_{x_{0},y_{0}}=0$ we considered the sign of the quantity
@@
which should be positive for a magnetic axis and negative for an X-point. For the up-down symmetric configurations considered, the magnetic axis is located at the same position, $(x=1,y=0)$ as the generalized Solov’ev configuration and $U=0$ thereon. Note that in this case the condition $dU/dy|_{(1,0)}=0$ is identically satisfied because of symmetry. If not otherwise specified, we assigned the following parametric values: $s=0.1$, $\delta=0.1$, $\lambda=0.1$ and $a_{0}=1$. It is recalled that when $\beta\neq 0$ and $\epsilon\neq 0$, $\delta$ should be calculated from (39). The parameters $C_{0}$, $C_{1}$ and $D_{1}$ were determined as solutions to the set of the algebraic equations derived by the conditions associated with the imposition of a stagnation point. Since we employed dimensionless quantities, the reference dimensional parameters $R_{0}$, $P_{s0}$ (in units of $4\pi\times 10^{-7}$ H/mPA ) and $U_{0}$ remained free. Depending on the value of $\epsilon$ we constructed tokamak, spherical tokamak and spheromak equilibria, the latter corresponding to $\epsilon=0$.
@@
First, we examined the impact of the $\alpha$- $\beta$- and $\gamma$-terms individually on up-down symmetric equilibria by imposing the same position of the magnetic axis, $(x=1,y=0)$, as that of the respective generalized Solov’ev equilibrium and the same value $U=0$ thereon. In the presence of $\alpha$-term, a D-shaped configuration is formed with negative triangularity, the boundary of which is shown in Fig. 2-left by the red-dashed curve. As $\alpha$ gradually takes larger values the configuration with increasing triangularity extends toward the axis of symmetry, eventually touching it to form a separatrix with a single X-point located at the origin of the coordinate system (blue, dot-dashed line) and then becomes a peculiar configuration having a separatrix with a couple of X-points (green-dotted line).
@@
The $\beta$-term creates configurations with small triangularity elongated along the $x$-axis as shown in Fig. 3. The red-dashed curves represent the magnetic surfaces of a configuration with the same poloidal magnetic flux as the respective generalized Solov’ev tokamak configuration represented by the black-continuous curve. As $\beta$ takes larger values the configuration extends further and eventually reaches the axis of symmetry transforming into a spheromak. It is noted that for $\epsilon=\beta=\gamma=0$ the equilibrium becomes independent of $\beta$. The impact of the $\gamma$-term is similar to that of the $\alpha$-term as shown in Fig. 4.
@@
Subsequently, the impact of combinations of the $\alpha$-, $\beta$- and $\gamma$-terms was examined by imposing an X-point at several positions. Thus, a variety of configurations were created. Three examples are given in Figs. 5-7. In Fig. 5 the blue, dot-dashed curves represent an up-down symmetric spherical tokamak configuration ($\epsilon=0.1$) the X-points of which are common with those of the respective generalized Solov’ev equilibrium, but the poloidal magnetic flux of this former equilibrium is larger than that of the latter. The red-dashed curve represents the boundary of a D-shaped configuration with poloidal magnetic flux equal to that of the generalized Solov’ev solution. In Fig. 6 is shown an up-down symmetric diverted configuration with the imposed X-points located at ($x=0.84,y=\pm 14.5$), while Fig. 7 shows an up-down asymmetric configuration with the magnetic axis located at ($x=1.02,y=0.83$) and the single lower X-point at $(x=0.94,y=-22.6)$.
@@
Figure 3: Fig. 3 shows the impact of the additional linear-in-$U$, poloidal-current term in the second of (2) in connection with different values of $\beta$. The black continuous curves represent the magnetic surfaces of the respective generalized Solov’ev tokamak configuration. $U_{b}$ is the value of the bounding $u$-curve (separatrix).
@@
Figure 4: Left: different equilibrium boundaries are given as the additional linear-in-$U$, electric-field term in the third of (2) varies in connection with different values of $\gamma$. The black continuous curve represents the separatrix of the respective generalized Solov’ev spheromak configuration. Right: A set of magnetic surfaces for one of the configuration represented by green-dotted U-curves are shown together with the respective curves of the Solov’ev equilibrium. $U_{b}$ is the value of the bounding $u$-curve (separatrix).
@@
Figure 5: A spherical tokamak diverted configuration ($\epsilon=0.1$, $\delta=20$) with a couple of X-points common with the respective Solov’ev equilibrium. The red-dashed curve represents the boundary of a D-shaped tokamak equilibrium with poloidal magnetic flux equal to that of the Solov’ev one. $U_{b}$ is the value of the bounding $u$-curve (separatrix).
@@
Figure 6: An up-down symmetric, diverted tokamak configuration ($\epsilon=0.1,\ \delta=20$) with a couple of X-points located at ($x=0.84,y=\pm 14.5$). The black continuous curves represent the magnetic surfaces of the respective Solov’ev equilibrium. $U_{b}$ is the value of the bounding $u$-curve (separatrix).
@@
Figure 7: An up-down asymmetric, diverted tokamak configuration for $k=0.2$, $\epsilon=0.2$, $\delta=31.6$, $D_{1}=0.5$ and $D_{2}=0.01$. The magnetic axis is located at $(x=1.02,y=0.83)$ and the single lower X-point at $(x=0.94,y=-22.6)$. The black continuous curves represent the magnetic surfaces of the respective Solov’ev equilibrium. $U_{b}$ is the value of the bounding $U$-curve (separatrix).
Finally, to completely determine the equilibrium two from the five free functions involved should be specified. For example, considering as free functions the density and the poloidal Mach function we make the choice
$\rho=\rho_{0}\left(1-\frac{U}{U_{b}}\right)^{l_{\rho}}\,,\ \ M_{p}=M_{p0}\left(1-\frac{U}{U_{b}}\right)^{l_{M_{p}}}\,,$ (40)
where $\rho_{0}$, $M_{p0}$, $l_{\rho}$ and $l_{M_{p}}$ are free parameters. Introducing the dimensionless quantities
$\tilde{\rho}=\frac{\rho}{\rho_{0}}\,,\ \ \tilde{X}=\frac{X}{u_{0}/R_{0}}\,,\ \ \tilde{{\mathbf{B}}}=\frac{{\mathbf{B}}}{u_{0}/R_{0}^{2}}\,,\ \ \tilde{{\mathbf{j}}}=\frac{{\mathbf{j}}}{u_{0}/(R_{0}^{3}\mu_{0})}\,,\ \ \tilde{{\mathbf{v}}}=\frac{{\mathbf{v}}}{u_{0}/(R_{0}^{2}\sqrt{\mu_{0}\rho_{0}})}\,,$
$\tilde{\Phi}=\frac{\Phi}{u_{0}^{2}/(R_{0}^{3}\sqrt{\mu_{0}\rho_{0}})}\,,\ \ \tilde{P}=\frac{P}{P_{s0}/(u_{0}^{2}/R_{0}^{4})}\,,$
integration of the relations (2) with respect to $u$ yields (in the dimensionless form and omitting the tildes)
$P_{s}$ $=$ $A-2(1+\delta^{2})U+\alpha(1+\delta^{2})U^{2}$
$\frac{X^{2}}{1-M_{p}^{2}}$ $=$ $B+4\epsilon U-2\beta U^{2}$
$\rho\left(\frac{d\Phi}{dU}\right)^{2}$ $=$ $C-4\lambda U+2\gamma U^{2}\,,$ (41)
where $A$, $B$ and $C$ are free parameters. The reference density $\rho_{0}$ remains free. Using (4) in (8) we obtain
$P=A-2(1+\delta^{2})U+\alpha(1+\delta^{2})U^{2}-\frac{1}{2}\rho v^{2}+x^{2}\left(C-4\lambda U+2\gamma U^{2}\right).$ (42)
The values of $A$ and $C$ are fixed so that the pressure vanish on the boundary. Isobaric contours for the D-shaped tokamak configuration of Fig. 5, the boundary of which is indicated by the red-dotted curve therein, are given in Fig. 8 together with pressure profiles on the poloidal plane at $y=0$ and at the vertical line passing from the position of the pressure maximum. It is noted that although the values of the dimensionless pressure are quite high, realistic pressure values can be obtained by appropriate values of the free parameters $P_{s0}$ and $U_{0}$. For flows of experimental fusion pertinence the deviation of the magnetic surfaces from the isobaric surfaces is very small.
Figure 8: Upper: Isobaric (blue-continuous) curves on the poloidal plane for the D-shaped tokamak configuration of Fig. (5) for $B=0.2$, $M_{p0}=0.1$, $l_{\rho}=2$ and $l_{M_{p}}=2$. The red-dashed curves represent magnetic surfaces. Lower left: Respective pressure profile on the mid-plane $y=0$. Lower right: Respective vertical pressure profile at the position of the pressure maximum
@@ 5 Conclusions
Making a generic linearizing choice of the free function-terms involved in the generalized GS equation (3) we solved the resulting linearized equation (2) analytically. The requested solution was expressed as an expansion of the generalized Solov’ev solution (12) in the form (15) involving the determinable function $W$. Then, the solution was obtained as a superposition of the counterpart homogeneous equation and a particular solution of the complete inhomogeneous equation in the form (28). The solutions of the pertinent radial ODEs were expressed in terms of infinite converging series.
Employing these solutions, we constructed several configurations by changing the values of the free parameters, either individually or in combination, associated with the new terms of pressure, poloidal current in conjunction with the parallel velocity, and electric field. The equilbria constructed are pertinent to tokamaks, spherical tokamaks and spheromaks; they include D-shaped configurations with either positive or negative triangularity and diverted configurations with either a couple of X-points or a single X-point.
@@
This work was conducted in the framework of participation of the University of Ioannina in the National Programme for the Controlled Thermonuclear Fusion, Hellenic Republic. The authors would like to thank the anonymous reviewers for constructive comments, which have resulted in a remarkably improved version of the manuscript.